OpenStudy (chelseabrown):

A line has a slope of 1/2 and a run = 50. What is the rise?

OpenStudy (loginusernamed):

25? (i dont now)

OpenStudy (google):

...

OpenStudy (loginusernamed):

:(

OpenStudy (mrnood):

slope = rise/run so in your case 1/2 = rise/50

OpenStudy (chelseabrown):

okay thought so... thank you guys!